How do I go about solving this question?
Hiking up to Chimney Pond on Mt. Katahdin takes 5 hours, but hiking down takes only 1 1/2 hours. The distance is 3.3 miles. What is the difference between the two average hiking rates?
Thanks in advance.

Answer by Theo(13342)

rate * time = distance.

5 hours to hike up.
1.5 hours to hike down.
distance is 3.3 miles.


going up, the formula becomes rate * 5 = 3.3 miles.

solve for rate to get rate = 3.3 / 5 = .66 miles per hour.

coming back down, the formula becomes rate * 1.5 = 3.3 miles.

solve for rate to get rate = 3.3 / 1.5 = 2.2 miles per hour.

the difference between the two average hiking rates is 2.2 minus .66 = 1.54 miles per hour.